Solve the following initial value problem: 4-3 )+sin Compute Select one: A. B. C. D. (2e1
Answer #1

4) (-)-(-sin(t)) d-/3 -sin)- given x =Ax+f(t),here consider the characteristic polynomial det(A-A1)- f (t)- 26 3 4-3- therefotherefore the general solution is 1/2 5t now the matrix X = (xi %) to findy, by using variation of parametres 5t 4 -sin (t) 0finally Xp = X1 X- Я--e: cos(t)+1 est sin (t) St ー-sin(t 13 13 65 general solution is x(r)eqre- -1/2) . | 26cos(t)-2-sin (r)